Given these balanced half-reactions

3H2O+I??IO3?+6H++6e?

Cl2+2e??2Cl?

enter the overall balanced redox reaction in an acidic solution.
Answer #1

3 H2O + I- = IO3- + 6 H+ + 6e-
Cl2 + 2 e- = 2 Cl-

3 H2O + I- + 3 Cl2 =IO3- + 6 H+ + 6 Cl-